If the density of a substance is 2.70g/L, what is the vol of 387 g of the substance?  

Note: I answered 143.33 L which was wrong
Answer #1

we Mhow that density (d) = mass (m) Vollemelv) from data d =2.709% mass = m = 3879 :: Vaht = 3879 2.709/ = 143.333L 1000ml =if d = 2.709/m1 ; m = 3879 Van - 3879 - 5 14 3:33 mi 2.709/mi Volume = 143.3 ml E 143 mi CScanned with CamScanner